# Markov spin models for image generation : explicit large deviations with respect to the number of pixels

### v1

@@ II.2 Explicit choice for the trivial forward noising dynamics
The forward dynamics should converge towards some simple steady state $P_{*}(\vec{S})$ representing ’noise’ that can be easily generated. The simplest choice is when the $N$ pixels evolve towards ’noise’ independently, i.e. when the forward propagator $\langle\vec{S}(t)|{\cal W}|\vec{S}(t-1)\rangle$ is factorized over the propagators $\langle S_{n}(t)|W|S_{n}(t-1)\rangle$ associated to the $N$ pixels $n=1,..,N$
@@
The simplest way to choose the $2\times 2$ Markov matrix $W$ amounts to use the orthonormalized eigenvectors associated to the two eigenvalues (±)plus-or-minus($\pm$)( ± ) of the Pauli matrix $\sigma^{x}$
@@
in order to write the spectral decomposition of $W$ involving the largest eigenvalue unity $\lambda_{0}=1$ and the other eigenvalue $\lambda\in]0,1[$ that parametrizes the characteristic relaxation time towards pure noise
@@
The physical interpretation is that during one-time-step, the spin $S$ will flip towards $(-S)$ with the probability $\frac{1-\lambda}{2}$ and will keep its value $S$ with the complementary probability $\frac{1+\lambda}{2}$.
The power $t$ of the matrix $W$ of Eq. 14
$W^{t}=|\psi_{0}\rangle\langle\psi_{0}|+\lambda^{t}|\psi_{1}\rangle\langle\psi_{1}|=\begin{pmatrix}\frac{1+\lambda_{1}^{t}}{2}&\frac{1-\lambda_{1}^{t}}{2}\\ \frac{1-\lambda_{1}^{t}}{2}&\frac{1+\lambda_{1}^{t}}{2}\end{pmatrix}=\frac{1+\lambda^{t}}{2}\mathbb{1}+\frac{1-\lambda^{t}}{2}\sigma^{x}$ (15)
@@
where the various expressions are based on the fact that each spin can only take the two values (±)plus-or-minus($\pm$)( ± ). These expressions show the essential role played by the overlap $Q$ between the image $\vec{S}$ at time $t$ and the image $\vec{\sigma}$ at time $t=0$
@@
with its two contributions $Q_{\pm}$ counting the number of pixels $n$ that have the same value or not in the two images
@@
Note that the overlap $Q$ of Eq. 19 has the same parity as $N$ and can take only the following $(N+1)$ values
@@
For large time $t\to+\infty$, the forward propagator of Eq. 18 converges for any initial condition towards the steady distribution
@@
where the $2^{N}$ configurations have all the same probability $\frac{1}{2^{N}}$.
@@ II.3 Direct link with the continuous-time Markov spin-flip dynamics of [14]
The finite-time propagator of Eq. 18 concerning the discrete-time Markov chain dynamics for the spins has exactly the same form as the finite-time propagator written in Eq. 83 of [14] for the continuous-time Markov spin-flip dynamics up to the correspondance of notations $\lambda=e^{-e_{1}}$. As a consequence, all the results that are based on this explicit finite-time propagator can be directly translated between the discrete-time and the continuous-time frameworks. While the work [14] focuses on the temporal convergence properties via the spectral decomposition of the forward propagator of Eq. 18 with its consequences for observables like the local magnetizations ${\cal M}_{i}(t)$ and the correlations ${\cal C}_{i_{1},i_{2},..,i_{K}}(t)$ between an arbitrary number $K$ of spins $1\leq i_{1}<i_{2}<..,i_{K}\leq N$
@@
the goal of the following sections is to analyze the full probability distributions of extensive global variables corresponding to sums over the $N$ pixels : we have already stressed that the overlap $Q$ of Eq. 19 is the relevant variable that determines the forward propagator of Eq. 18, but we will also consider the magnetization
@@
in order to extract the large deviations properties of the corresponding intensive variables $q=\frac{Q}{N}$ and/or $m=\frac{M}{N}$ when the number $N$ of pixels becomes large $N\to+\infty$ during the forward dynamics and during the backward-generative dynamics.
@@ III Large deviations properties of the forward propagator for $N$ pixels
In this section, we analyze the large deviations properties of the forward propagator $\langle\vec{S}|{\cal W}^{t}|\vec{\sigma}\rangle$ of Eq. 18 when the number $N$ of pixels is large $N\to+\infty$.
III.1 Joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ of the magnetization $M$ and of the overlap $Q$ with the initial condition $\vec{\sigma}$
III.1.1 Explicit joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ for any finite size $N$
For a given initial condition $\vec{\sigma}$, the joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ at time $t$ of the magnetization $M$ and of the overlap $Q$ produced by the forward propagator $\vec{S}|{\cal W}^{t}|\vec{\sigma}\rangle$ of Eq. 18 reads
@@ III.1.1 Explicit joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ for any finite size $N$
$=\sum_{\vec{S}}\delta_{M,\sum_{n=1}^{N}S_{n}}\delta_{Q,\sum_{n=1}^{N}S_{n}\sigma_{n}}\left(\frac{1+\lambda^{t}}{2}\right)^{\frac{N+Q}{2}}\left(\frac{1-\lambda^{t}}{2}\right)^{\frac{N-Q}{2}}$
@@
determines the corresponding populations of up and down spins
@@
In order to analyze the joint properties with $M$ and $Q$, one needs to introduce the four populations labelled by ϵ=±italic-$\epsilon=\pm$ and ϵ0=±subscriptitalic-$\epsilon_{0}=\pm$
@@
As a consequence, the joint distribution of Eq. 25 depends on the initial condition $\vec{\sigma}$ only via its magnetization $M_{0}$ and involves the two binomial coefficients associated to the combinatorics of the two decompositions of Eqs 30 and 31
@@
III.1.2 Large deviations for large $N$ : rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ for the intensive magnetization $m$ and the intensive overlap $q$
@@
The Stirling formula for the logarithm of the factorial
leads to the contributions
$\sum_{\eta=\pm 1}\ln\left[\left(N\frac{1+\eta m_{0}}{2}\right)!\right]$
and
$\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\ln\left[N\left(\frac{1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q}{4}\right)!\right]$
so that their difference needed for Eq. 33 reduces to
$\sum_{\eta=\pm 1}\ln\left[\left(N\frac{1+\eta m_{0}}{2}\right)!\right]-\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\ln\left[N\left(\frac{1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q}{4}\right)!\right]$
The conclusion is that Eq. 33 displays the following large deviation behavior for large $N$
@@
where the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ reads
@@
It is convenient to keep the last term as a sum over the four terms $(\epsilon=\pm 1;\epsilon_{0}=\pm 1)$ in order to simplify the computations of the first-order partial derivatives with respect to $m$
$\partial_{m}{\cal I}_{t}^{[m_{0}]}(m,q)$ $=\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{\epsilon}{4}\ln\left(1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q\right)$ (41)
$=\frac{1}{4}\ln\left(1+m+m_{0}+q\right)+\frac{1}{4}\ln\left(1+m-m_{0}-q\right)-\frac{1}{4}\ln\left(1-m+m_{0}-q\right)-\frac{1}{4}\ln\left(1-m-m_{0}+q\right)$
$=\frac{1}{4}\ln\left(\frac{(1+m)^{2}-(m_{0}+q)^{2}}{(1-m)^{2}-(m_{0}-q)^{2}}\right)$
and with respect to $q$
$\partial_{q}{\cal I}_{t}^{[m_{0}]}(m,q)=-\frac{1}{2}\ln\left(\frac{1+\lambda^{t}}{2}\right)+\frac{1}{2}\ln\left(\frac{1-\lambda^{t}}{2}\right)+\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{\epsilon\epsilon_{0}}{4}\ln\left(1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q\right)$
$=\frac{1}{2}\ln\left(\frac{1-\lambda^{t}}{1+\lambda^{t}}\right)+\frac{1}{4}\ln\left(1+m+m_{0}+q\right)+\frac{1}{4}\ln\left(1-m-m_{0}+q\right)-\frac{1}{4}\ln\left(1+m-m_{0}-q\right)-\frac{1}{4}\ln\left(1-m+m_{0}-q\right)$
$=-\frac{1}{2}\ln\left(\frac{1+\lambda^{t}}{1-\lambda^{t}}\right)+\frac{1}{4}\ln\left(\frac{(1+q)^{2}-(m+m_{0})^{2}}{(1-q)^{2}-(m-m_{0})^{2}}\right)$ (42)
and of the second-order partial derivatives
$\partial_{m}^{2}{\cal I}_{t}^{[m_{0}]}(m,q)$ $=\frac{1}{4}\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{1}{1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q}=\partial_{q}^{2}{\cal I}_{t}^{[m_{0}]}(m,q)$ (43)
@@
One can use the first derivatives of Eqs 41 42 to rewrite the rate function of Eq. 40 as
At the typical value $(\hat{m},\hat{q})$ of the pair $(m,q)$, the vanishing of the first derivatives of Eqs 41 and 42
@@
where the rate function of Eq. 45 also vanishes
as it should.
It is thus convenient to use this equation to replace
whose vanishing at the typical value $(\hat{m},\hat{q})$ of the pair $(m,q)$ is now obvious.
Note that the Taylor series of the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ at second order around the typical value $(\hat{m},\hat{q})$ of the pair $(m,q)$ that involves the second derivatives of Eqs 43 and 44
${\cal I}_{t}^{[m_{0}]}(m,q)$ $=\frac{(m-\lambda^{t}m_{0})^{2}+(q-\lambda^{t})^{2}}{2}\left[\frac{1}{4}\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{1}{1+\epsilon\lambda^{t}m_{0}+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}\lambda^{t}}\right]$ (51)
$+(m-\lambda^{t}m_{0})(q-\lambda^{t})\left[\frac{1}{4}\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{\epsilon_{0}}{1+\epsilon\lambda^{t}m_{0}+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}\lambda^{t}}\right]+...$
governs the small Gaussian fluctuations of the pair $(m,q)$ around the typical value $(\hat{m},\hat{q})$.
III.2 Generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ of the magnetization $M$ and of the overlap $Q$ with the initial condition $\vec{\sigma}$
III.2.1 Generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ for any finite $N$
Instead of the joint probability distribution $P^{[\vec{\sigma}]}_{t}(M,Q)$ of Eq. 25, one can focus on the corresponding generating function

### v2

@@ II.2 Explicit choice for the trivial forward noising dynamics
Since the forward dynamics should converge towards some simple steady state $P_{*}(\vec{S})$ representing ’noise’ that can be easily generated, the simplest choice is when the $N$ pixels evolve towards ’noise’ independently, i.e. when the forward propagator $\langle\vec{S}(t)|{\cal W}|\vec{S}(t-1)\rangle$ is factorized over the propagators $\langle S_{n}(t)|W|S_{n}(t-1)\rangle$ associated to the $N$ pixels $n=1,..,N$
@@
Note that in the field of stochastic dynamics of many-body spin or particle systems where the main goal is to analyze the dynamical consequences of the interactions, this choice of a non-interacting dynamics is actually never considered and would appear as too trivial to be interesting on its own, while from the point of view of generative Markov models, this choice of a trivial forward dynamics is fully justified by the further purpose of constructing the non-trivial backward dynamics where the forward propagator plays the role of a simple building block, while the non-trivial part enters via the initial distribution $P^{ini}$ of the interesting images.
@@
The simplest way to choose the $2\times 2$ Markov matrix $W$ concerning a single pixel $S=\pm$ amounts to use the well-known orthonormalized eigenvectors associated to the two eigenvalues (±)plus-or-minus($\pm$)( ± ) of the Pauli matrix $\sigma^{x}$
@@
in order to write the spectral decomposition of $W$ involving the largest eigenvalue unity $\lambda_{0}=1$ that is required by the probability conservation and the other smaller eigenvalue $\lambda\in]0,1[$ that will parametrize the characteristic relaxation time towards pure noise
@@
The physical interpretation is that during one-time-step, the spin $S$ will flip towards $(-S)$ with the probability $\frac{1-\lambda}{2}$ and will keep its value $S$ with the complementary probability $\frac{1+\lambda}{2}$
$\langle+|W|-\rangle=\langle-|W|+\rangle$ $=\frac{1-\lambda}{2}$ (15)
$\langle+|W|+\rangle=\langle-|W|-\rangle$ $=\frac{1+\lambda}{2}$ (16)
The power $t$ of the Markov matrix $W$ of Eq. 14
$W^{t}=|\psi_{0}\rangle\langle\psi_{0}|+\lambda^{t}|\psi_{1}\rangle\langle\psi_{1}|=\begin{pmatrix}\frac{1+\lambda^{t}}{2}&\frac{1-\lambda^{t}}{2}\\ \frac{1-\lambda^{t}}{2}&\frac{1+\lambda^{t}}{2}\end{pmatrix}=\frac{1+\lambda^{t}}{2}\mathbb{1}+\frac{1-\lambda^{t}}{2}\sigma^{x}$ (17)
@@
The equivalence between the three alternative expressions is based on the fact that each spin can only take the two values (±)plus-or-minus($\pm$)( ± ). The two last expressions show the key role played by the extensive overlap $Q$ between the image $\vec{S}$ at time $t$ and the image $\vec{\sigma}$ at time $t=0$
@@
with the two contributions $Q_{\pm}$ counting the numbers of pixels $n$ that have the same value or not in the two images
@@
Note that the overlap $Q$ of Eq. 21 has the same parity as the number $N$ of pixels and can take only the following $(N+1)$ values
@@
For large time $t\to+\infty$, the forward propagator of Eq. 20 converges for any initial condition $\vec{\sigma}$ towards the steady distribution
@@
where the possible $2^{N}$ configurations have all the same probability $\frac{1}{2^{N}}$.
@@ II.3 Direct link with the continuous-time Markov spin-flip dynamics of [16]
The finite-time propagator of Eq. 20 obtained above for the discrete-time Markov chain dynamics has exactly the same form as the finite-time propagator written in Eq. 89 of [16] for the continuous-time Markov spin-flip dynamics up to the correspondance of notations $\lambda=e^{-e_{1}}$. As a consequence, all the results that are based on this explicit finite-time propagator can be directly translated between the discrete-time and the continuous-time frameworks. While the work [16] focuses on the temporal convergence properties via the spectral decomposition of the forward propagator of Eq. 20 with its consequences for observables like the local magnetizations ${\cal M}_{i}(t)$ and the correlations ${\cal C}_{i_{1},i_{2},..,i_{K}}(t)$ between an arbitrary number $K$ of spins $1\leq i_{1}<i_{2}<..,i_{K}\leq N$
@@
the goal of the following section will be to analyze the full probability distributions of extensive global variables corresponding to sums over the $N$ pixels: we have already stressed that the overlap $Q$ of Eq. 21 is the relevant variable that determines the forward propagator of Eq. 20, but we will also consider the global magnetization
@@
that represents the simplest extensive variable associated to a single configuration. When the number $N$ of pixels becomes large $N\to+\infty$, we will be interested into the large deviations properties of the corresponding intensive variables $q=\frac{Q}{N}$ and/or $m=\frac{M}{N}$ during the forward dynamics and into the consequences for the backward-generative dynamics.
@@ III Large deviations properties of the forward propagator for $N$ pixels
In this section, the goal is to analyze the statistical properties of the forward propagator $\langle\vec{S}|{\cal W}^{t}|\vec{\sigma}\rangle$ of Eq. 20 for any finite number $N$ of pixels and to extract the large deviations properties in the limit $N\to+\infty$.
III.1 Distribution $P_{t}(Q)$ of the overlap $Q$ with the initial condition
The probability distribution $P_{t}(Q)$ of the overlap $Q$ at time $t$ produced by the forward propagator $\langle\vec{S}|{\cal W}^{t}|\vec{\sigma}\rangle$ of Eq. 20
$P_{t}(Q)$ $\equiv\sum_{\vec{S}}\delta_{Q,\sum_{n=1}^{N}S_{n}\sigma_{n}}\langle\vec{S}|{\cal W}^{t}|\vec{\sigma}\rangle=\left(\frac{1+\lambda^{t}}{2}\right)^{\frac{N+Q}{2}}\left(\frac{1-\lambda^{t}}{2}\right)^{\frac{N-Q}{2}}\sum_{\vec{S}}\delta_{Q,\sum_{n=1}^{N}S_{n}\sigma_{n}}$ (27)
$=\left(\frac{1+\lambda^{t}}{2}\right)^{\frac{N+Q}{2}}\left(\frac{1-\lambda^{t}}{2}\right)^{\frac{N-Q}{2}}\frac{N!}{\frac{N+Q}{2}!\frac{N-Q}{2}!}$
reduces to the binomial distribution.
For large $N$, the Stirling approximation for the logarithm of the factorial
leads to the following large deviation behavior the intensive overlap $q=\frac{Q}{N}$
$P_{t}(Q=Nq)\mathop{\propto}_{N\to+\infty}e^{-Ni_{t}(q)}$ (29)
with the standard rate function $i_{t}(q)$ produced by the binomial distribution
$i_{t}(q)=\frac{1+q}{2}\ln\left(\frac{1+q}{1+\lambda^{t}}\right)+\frac{1-q}{2}\ln\left(\frac{1-q}{1-\lambda^{t}}\right)$ (30)
At the typical value ${\hat{q}}=\lambda^{t}$, the rate function $i_{t}(q)$ vanishes $i_{t}({\hat{q}})=0$ together with its first derivative
In the following subsection, the goal is to generalize this simple analysis to take into account the magnetization.
III.2 Joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ of the magnetization $M$ and of the overlap $Q$ with the initial condition $\vec{\sigma}$
III.2.1 Explicit joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ for any finite number $N$ of pixels
For a given initial condition $\vec{\sigma}$, the joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ of the magnetization $M$ and of the overlap $Q$ produced by the forward propagator $\vec{S}|{\cal W}^{t}|\vec{\sigma}\rangle$ at time $t$ of Eq. 20 reads
@@ III.2.1 Explicit joint distribution $P_{t}^{[\vec{\sigma}]}(M,Q)$ for any finite number $N$ of pixels
$=\left(\frac{1+\lambda^{t}}{2}\right)^{\frac{N+Q}{2}}\left(\frac{1-\lambda^{t}}{2}\right)^{\frac{N-Q}{2}}\sum_{\vec{S}}\delta_{M,\sum_{n=1}^{N}S_{n}}\delta_{Q,\sum_{n=1}^{N}S_{n}\sigma_{n}}$
@@
determines the corresponding populations of up and down spins in the initial condition
@@
In order to analyze the joint properties of $M$ and $Q$, one needs to introduce the four populations labelled by ϵ=±italic-$\epsilon=\pm$ and ϵ0=±subscriptitalic-$\epsilon_{0}=\pm$
@@
As a consequence, the joint distribution of Eq. 32 depends on the initial condition $\vec{\sigma}$ only via its magnetization $M_{0}$ and involves the two binomial coefficients associated to the combinatorics of the two decompositions of Eqs 37 and 38
@@
III.2.2 Large deviations for large $N$ : rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ for the intensive magnetization $m$ and the intensive overlap $q$
@@
As described in more details in the Appendix section A.1, Eq. 40 displays the following large deviation behavior for large $N$
@@
with the rate function
@@
Let us compute the first-order partial derivatives with respect to the magnetization $m$
$\partial_{m}{\cal I}_{t}^{[m_{0}]}(m,q)$ $=\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{\epsilon}{4}\ln\left(1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q\right)=\frac{1}{4}\ln\left(\frac{(1+m)^{2}-(m_{0}+q)^{2}}{(1-m)^{2}-(m_{0}-q)^{2}}\right)$ (44)
and with respect to the overlap $q$
$\partial_{q}{\cal I}_{t}^{[m_{0}]}(m,q)$ $=-\frac{1}{2}\ln\left(\frac{1+\lambda^{t}}{2}\right)+\frac{1}{2}\ln\left(\frac{1-\lambda^{t}}{2}\right)+\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{\epsilon\epsilon_{0}}{4}\ln\left(1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q\right)$ (45)
$=-\frac{1}{2}\ln\left(\frac{1+\lambda^{t}}{1-\lambda^{t}}\right)+\frac{1}{4}\ln\left(\frac{(1+q)^{2}-(m+m_{0})^{2}}{(1-q)^{2}-(m-m_{0})^{2}}\right)$
as well as the second-order partial derivatives
$\partial_{m}^{2}{\cal I}_{t}^{[m_{0}]}(m,q)=\partial_{q}^{2}{\cal I}_{t}^{[m_{0}]}(m,q)$ $=\frac{1}{4}\sum_{\epsilon=\pm 1}\sum_{\epsilon_{0}=\pm 1}\frac{1}{1+\epsilon m+\epsilon_{0}m_{0}+\epsilon\epsilon_{0}q}$ (46)
$=\frac{1}{2}\left[\frac{1+m_{0}}{(1+m_{0})^{2}-(m+q)^{2}}+\frac{1-m_{0}}{(1-m_{0})^{2}-(m-q)^{2}}\right]$
@@
$=\frac{1}{2}\left[\frac{1+m_{0}}{(1+m_{0})^{2}-(m+q)^{2}}-\frac{1-m_{0}}{(1-m_{0})^{2}-(m-q)^{2}}\right]$
At the typical value $(\hat{m},\hat{q})$ of the pair $(m,q)$, the vanishing of the first derivatives of Eqs 44 and 45
@@
where the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ itself also vanishes
${\cal I}_{t}^{[m_{0}]}(\hat{m}=\lambda^{t}m_{0},\hat{q}=\lambda^{t})=0$ (50)
The second derivatives of Eqs 46 and 47 evaluated for $(m,q)=(\hat{m},\hat{q})$
$\partial_{m}^{2}{\cal I}_{t}^{[m_{0}]}(m,q)\big{|}_{(m,q)=(\hat{m},\hat{q})}=\partial_{q}^{2}{\cal I}_{t}^{[m_{0}]}(m,q)\big{|}_{(m,q)=(\hat{m},\hat{q})}$ $=\frac{1}{(1-m_{0}^{2})(1-\lambda^{2t})}$
$\partial_{q}\partial_{m}{\cal I}_{t}^{[m_{0}]}(m,q)\big{|}_{(m,q)=(\hat{m},\hat{q})}$ $=-\frac{m_{0}}{(1-m_{0}^{2})(1-\lambda^{2t})}$ (51)
lead to the following Taylor series of the rate function ${\cal I}_{t}^{[m_{0}]}(m,q)$ at second order around the typical value $(\hat{m}=\lambda^{t}m_{0},\hat{q}=\lambda^{t})$ of the pair $(m,q)$ where the function and the two first derivatives vanish (Eqs 48 and 50)
${\cal I}_{t}^{[m_{0}]}(m,q)=\frac{(m-\lambda^{t}m_{0})^{2}+(q-\lambda^{t})^{2}-2m_{0}(m-\lambda^{t}m_{0})(q-\lambda^{t})}{2(1-m_{0}^{2})(1-\lambda^{2t})}+...$ (52)
that governs the joint Gaussian typical fluctuations of order $\frac{1}{\sqrt{N}}$ for the pair $(m,q)$ around the typical value $(\hat{m}=\lambda^{t}m_{0},\hat{q}=\lambda^{t})$
$p^{[m_{0}]}_{t}(m,q)\mathop{\propto}_{N\to+\infty}e^{\displaystyle-N{\cal I}_{t}^{[m_{0}]}(m,q)}\mathop{\propto}_{N\to+\infty}e^{\displaystyle-N\frac{(m-\lambda^{t}m_{0})^{2}+(q-\lambda^{t})^{2}-2m_{0}(m-\lambda^{t}m_{0})(q-\lambda^{t})}{2(1-m_{0}^{2})(1-\lambda^{2t})}+...}$ (53)
The prefactor of this bivariate Gaussian distribution can be reconstructed from the normalization condition when one integrates over the two variables $(m,q)$ in order to obtain the following Gaussian approximation for large $N$
$p^{[m_{0}]Gauss}_{t}(m,q)=\frac{1}{2\pi(1-\lambda^{2t})(1-m_{0}^{2})}e^{\displaystyle-N\frac{(m-\lambda^{t}m_{0})^{2}+(q-\lambda^{t})^{2}-2m_{0}(m-\lambda^{t}m_{0})(q-\lambda^{t})}{2(1-m_{0}^{2})(1-\lambda^{2t})}}$ (54)
III.3 Generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ of the magnetization $M$ and of the overlap $Q$ with the initial condition $\vec{\sigma}$
III.3.1 Generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ for any finite $N$
Instead of the joint probability distribution $P^{[\vec{\sigma}]}_{t}(M,Q)$ of Eq. 32, one can alternatively focus on the corresponding generating function ${\cal G}^{[\vec{\sigma}]}_{t}(h,g)$ involving the conjugated parameters $(h,g)$